An aluminium meter rod of area of cross section 4square cm with k=0.5cal per gm degrees centigrade is observed that at steady state 360cal of heat flows per minute.the tempareture gradiant along the rod is
 

Grade:12th pass

1 Answers

K Induvarshita
5 years ago
Given that A=4cmsquare. K=0.5cal/g°c,t= 60sec Q=360. we know that Q/t= KA×t1-t2/l.thenif we apply it360/60=0.5×4×t1-t2/l . then 6/2=t1-t2 that =3.